# Human-In-The-Loop Machine Learning for Safe and Ethical Autonomous Vehicles: Principles, Challenges, and Opportunities

### v1

@@ II-B1 Navigation and Data Collection
A salient challenge in UAV swarm is collision avoidance where UAVs must avoid other UAVs, fixed and moving object which necessiates arranging the input. The authors of [44] study the flocking and obstacle avoidance problem for multiple fixed-wing UAVs in cluttered environments. A novel curriculum-based multi-agent DRL is developed to learn the decentralized flocking and obstacle avoidance policy. In this context, the collision-avoiding flocking task is decomposed into multiple subtasks and the number of subtasks to be solved is gradually increased. In the developed approach, a novel multi-agent actor-critic DRL algorithm and two knowledge transfer mechanisms are developed for online learning and offline transfer, respectively. Similarly, the authors of [48] study collision-free flocking control problems for scalable and large-scale fixed-wing UAV swarms by integrating CL with multi-agent DRL. The problem considers both collision avoidance and the large population of UAVs in the swarm. The key idea is to divide the training process into several learning phases and gradually expand the scale of UAV swarms. In addition, the authors of [51] investigate navigating UAVs in complex 3D environments and develop a solution using curriculum-based DRL. The human expert divides the training into ten different levels with an increasing range of possible numbers of obstacles in the environment. The leveraged curriculum exposes the UAV to simpler environments initially and progressively to more complex ones.
@@
Moreover, data collection and UAV landing tasks require arranging the input. The authors of [52] address the collection of data from wireless sensor nodes to minimize the Age of Information during data collection, while considering the energetic sustainability of UAVs. They introduce Multi-Agent Reinforcement Learning (MARL) in combination with CL and an evolutionary strategy to improve the performance of traditional MARL setups as the number of agents increases. Their method promotes MARL scalability and integrates energy-efficient charging mechanisms that effectively improve system performance in large-scale deployments. Moreover, the collaborative autonomous landing of a quadrotor UAV on a moving Unmanned Ground Vehicle (UGV) is challenging due to the need for accurate real-time tracking of the UGV and the adjustment of the landing policy. The authors of [53] propose the Landing Vision System to offer rapid localization and pose estimation of the UGV. Then, they design an Automatic CL approach in conjunction with the Twin Delayed Deterministic Policy Gradient (TD3) that breaks down UAV landing control into a series of simpler subtasks, and learn the landing tasks under different conditions of UGV motions and wind interference. Table III summarizes the above findings.
II-B2 AD
@@
Self-driving tasks such as intersection management and motion planning require CL to faciliate the learning tasks. The authors of [16] present a reward-driven automated curriculum proximal policy optimization (RD-ACPPO) approach for self-driving tasks at unsignalized intersections. Simply, an adaptive CL technique is presented to accommodate a progressively increasing difficulty of learning tasks in AD. Their approach allows for the reasonable allocation of the importance weights for different curricula during various RL training periods. The RL policy trained by the proposed framework demonstrated reasonable curriculum transition timings during the training phase and attained the highest task success rate. In addition, the authors of [56] develop a novel CL scheme for motion planning in AD that incorporates traffic rule and map information into a GNN-RL agent whose state is extended by map and traffic rule information. They decompose the complicated learning task by applying CL for GNN-RL agent. The curriculum’s sequence of sub-tasks gradually increases the level of difficulty of the motion planning problem for the agent. Each of their sub-tasks contains a set of rewards, including new rewards for temporal-logic-based traffic rules for speed, safety distance, and braking. Moreover, the authors of [45] investigate the challenge of learning AD behaviors at urban intersections with DRL. They combine CL with DRL to reduce the required training time and improve agent performance. They formed a set of tasks to consider and used an automated method to create the curriculum for the training process to avoid the burden of human handcrafting. Table IV summarizes the above findings.
@@
CL can faciliate learning and training in CARLA driving simulator and target search for UAV swarm. The authors of [49] aim to learn an end-to-end urban driving policy for the CARLA driving simulator by combining the PPO algorithm with CL. In their approach, the RL phase is divided into several stages of increasing difficulty, so that the agent is guided to learn a better driving policy. The curriculum-based agent is evaluated on different metrics, cities, weather conditions, and traffic scenarios. Furthermore, the authors of [55] develop a multi-agent, scenario-based AD training and evaluation framework for the CARLA simulator. The results demonstrate the proposed approach is practical for multi-agent learning and can generate auto-curricula from scenario descriptions. In a similar vein, the authors of [54] use UAV swarm for target search in disaster management scenarios. An adaptive embedded curriculum can adjust the difficulty of the task based on the success rate achieved in training and is used to enable UAVs to explore and navigate cluttered 3D environments with sparse rewards efficiently.
@@
By systematically increasing the complexity of tasks, CL not only lays a solid foundation for the AVs’ learning process but also reduces training time and improves agent performance. This structured approach is crucial for preparing the vehicle to handle complex scenarios, where HITL-RL becomes necessary to fine-tune the vehicle’s decision-making abilities.
II-C Key Findings and Insights
In CL, human experts systematically train models by starting with simple tasks and gradually working their way up to more difficult tasks. Complex tasks in AVs can be broken down into smaller, more manageable subtasks. CL helps to organize these tasks from simple to complex, facilitating the learning process. This approach ensures optimal learning outcomes and faster convergence speed. Combining CL with other techniques such as RL can significantly improve the training process. In the context of AVs, this combination can reduce training time, improve agent performance, and facilitate the learning of complex tasks such as navigation, obstacle avoidance, and collision-free flocking control.
CL can also be applied to scalable and large-scale systems, such as large populations of UAV swarms. An adaptive embedded curriculum for UAV swarms adjusts task difficulty based on training success rates, enabling effective exploration and navigation in cluttered environments with sparse rewards. Also, CL with multi-agent DRL is scalable to large UAV swarms, allowing the gradual increase of swarm size and ensuring collision-free flocking. Further, automatic CL with TD3 for UAV landing on UGVs offers rapid localization and pose estimation, handles various UGV motions and wind interference, and breaks down tasks into simpler subtasks.
In the context of AD, the combination of CL with DRL at urban intersections shortens training time and improves agent performance by automating the creation of training plans and eliminating the burden of human manual labor. Meanwhile, dividing the RL phase into stages of increasing difficulty for End-to-End Urban Driving helps the agent learn better policies by guiding them through increasingly difficult tasks. Further, we learn that an adaptive CL technique enables a progressive increase in task difficulty, facilitating efficient learning and achieving high success rates in AD. Moreover, incorporating map and traffic rule information into the learning process helps the agent to master more complex motion planning tasks by progressively increasing the difficulty of tasks and providing specific rewards for rule-based behaviors. Overall, by dividing the training process into phases and gradually increasing the scope, CL can effectively train agents to cope with complex scenarios. Overall, integrating CL with AVs demonstrates its effectiveness in guiding autonomous agents through progressively challenging tasks, ultimately leading to improved performance and adaptability in real-world scenarios.
II-C1 Merger with Transfer Learning
@@
II-C2 CL VS Hard Example Mining
When developing a curriculum for AVs, a common approach is to sort tasks from easy to hard, which is central to CL. However, this raises the question of whether this incremental approach is always the most effective strategy. An alternative approach, Hard Example Mining (HEM)[58] , shifts the focus from a structured progression to selecting the most challenging examples for training. HEM operates under the assumption that harder examples, which are often identified by the model’s losses or gradient magnitude, provide more informative and valuable learning opportunities.
II-C3 Human Defined CL VS Human Data
@@
II-C4 Co-learning: a multi-agent approach to CL
@@ III Human-In-The-Loop Reinforcement Learning
This section reviews the state-of-the-art in HITL-RL and explores different methods through which human involvement can significantly enhance the RL process, leveraging techniques like reward shaping, directly injecting actions into the learning loop, and interactive learning, thereby assisting AVs in better fulfilling their mission, e.g., avoiding obstacles, navigating, and enhancing performance and safety along with fostering public trust.
Table V: Key HITL-DRL Techniques for UAV and AV Navigation
@@
III-A Preliminaries
We explore the state-of-the-art in HITL-RL and its application to AVs. HITL-RL significantly enhances the RL process by leveraging techniques like reward shaping, action injection, and interactive learning. These methods assist AVs in tasks such as lane-change, driving, obstacle avoidance, navigation, and overall performance enhancement, thereby improving safety and fostering public trust. The setup of HITL-RL for AVs is depicted in Fig. 4. In this setup, human experts play a crucial role in monitoring the RL process. They can intervene by injecting actions when necessary and can enhance the system by providing supplementary rewards, known as human rewards. This integration ensures that the learning process remains safe, reliable, and aligned with human values and judgments.
III-B Literature review
III-B1 Reward Shaping
In reward shaping, the environmental reward is supplemented with human-defined rewards to guide learning. HITL can switch between multiple UAV reward functions in different situations and assist in constructing interactive reward functions.
Human can help to switch between different reward in different situations and replace the manual construction of the reward function. The authors of [36] develop a HITL-DRL framework for UAV navigation in large-scale 3D complex environments.
This work proposes a DRL-based method combined with HITL, which allows the UAV to avoid obstacles automatically during flight. The role of HITL is to assist the UAV in obstacle avoidance by dynamically switching between multiple UAV reward functions in different situations. The combined approach significantly improves the success rate and navigation efficiency. In a similar vein, the authors of [59] address the problem of autonomous UAV navigation and develop a practical solution based on Preference-based Reinforcement Learning (PbRL). In PbRL, the manual construction of the reward function is replaced with human feedback. A human is presented with a pair of trajectories followed by the RL agent and is asked to indicate a preference for one over the other. The PbRL algorithm then computes an optimal sequence of actions using only the set of preferences collected over different trajectories.
In the AD context, reward can be learned from human driving data and encoded with human prior experience. The authors of [60] address the intricate challenges in designing appropriate reward functions to manage optimizing complex goals in RL. They develop a framework that learns the reward function from human driving data to help manage the learning process. Trajectory sample, offline preference learning, and RL are the comprising components, where feasible trajectories are generated by sampling end targets from a reachable state space, then a transformer network that is used to model the human driving decision-making process, and obtaining a reward function, is trained by a new offline preference learning where generated feasible trajectories are compared with human driving trajectories. Finally, the final driving decision network is obtained by incorporating the derived reward function in the RL framework. In another related effort, the authors of [61] aim to facilitate AVs mandatory lane-change scenarios in complex environments. In this context, they develop a HITL-RL method based on Twin Delayed Deep Deterministic policy gradient (TD3) for changing lane decisions. They encode human prior experience into reward function and safety constraints in an offline manner. Also, human experts guide the agents online to limit unsafe exploration during the training process and provide a demonstration in complex scenarios. Table V summarizes the above findings.
III-B2 Action Injection
@@
Introducing human guidance into the training loop of DRL algorithms enable AV to better handle different situations. Further, human intervention can make safe the UAV swarm. The authors of [37] state that the limited intelligence and capabilities of DRL-based strategies in AV applications prevent them from handling various situations and replacing human drivers. Introducing human guidance into the training loop of DRL algorithms can improve the situation. In the proposed HITL-DRL, human guidance is used in real-time to improve the agent’s learning ability and performance. Specifically, the human participant observes the agent’s real-time training process and assists when necessary. The provided action of the human guidance is prioritized as the output action of the agent to interact with the environment and replaces the original action of the DRL policy. The actor and critic are updated according to the human guidance and experience of the DRL agent. The actor-network can learn from both human guidance through imitation learning and the experience of the interactions through RL. The critic network can evaluate the values of the agent’s actions as well as the actions of human guidance. Similarly, the authors of [62] develop a HITL-DRL framework for UAV swarm navigation in a large-scale 3D complex urban environment. This framework enables human intervention in the flight of the UAV swarm when necessary to correct dangerous actions.
Table VI: Key HITL-DRL Techniques for UAV and AV Navigation
@@
Additionally, the Actor-Critic structure of DDPG has been modified to identify and learn from positive human experiences.
@@
Human feedback can enhance UAV obstacle avoidance performance. Further, human can act as as a mentor to the AI agent of AD. The authors of [63] investigate autonomous navigation in challenging environments with the capability to avoid obstacles. The UAV navigates toward its target using the DDPG model, which can handle unexpected situations. HITL feedback is integrated with the DDPG model to enhance obstacle avoidance performance. This feedback is crucial when the pre-trained DDPG model encounters environments it is not fully adapted to, and gives the system the flexibility to handle unforeseen scenarios effectively and make real-time adjustments. For the HCI feedback, the first-person shooter concept is employed, where feedback is provided using keyboard directional keys. Through this HCI-based HITL feedback, real-time control of the UAV’s flight movement is achieved. Moreover, the authors of [64] aim to improve the safety of AD technology, ensuring traffic flow efficiency in mixed-traffic environments. They develop an enhanced HITL-RL for driving policy learning, where the human serves as a mentor to the AI agent, supervising, intervening, and demonstrating in its learning process, and integrated with DRL to guarantee the safety of the AVs, but also optimize the traffic flow efficiency. Table VI summarizes the above findings.
III-B3 Interactive Learning
@@
The agent and human trainer can interact and cooperate to solve tasks. For example, human can supplement environmental reward and guide learning objectives to improve learning efficiency and RL performance. The authors of [38] aim to explore the oceans using autonomous underwater vehicles (AUV). RL can improve its autonomy but still has practical problems due to sparse rewards and low learning efficiency. They propose a deep interactive RL method for AUV path tracking by combining the advantages of DRL and interactive RL. Since the human trainer cannot give human rewards to the AUV in the operational condition and the AUV needs to adapt to a changing environment when it is in the ocean, a DRL method that simultaneously learns from human rewards and environmental rewards is proposed. The authors of [65] improve RL performance in simulation and practice for the navigation task of UGVs using HITL-RL. The framework improves human interventions with a series of mechanisms to provide demonstrations and improve RL’s capabilities. The mechanisms comprise human-guided learning objectives, prioritized human experience replay, and human intervention-based reward shaping.
@@
Human behavior modeling mechanisms can alleviate the workload on humans, further using meta-model provides a language to describe the roles of UAVs and humans and autonomous decisions. The authors of [40] aim to improve the learning performance of RL by incorporating human guidance. In this regard, they develop a human guidance-based RL framework along with a novel mechanism for leveraging the experience of human guidance. A TD3-based algorithm was developed to improve algorithmic capabilities in the context of HITL-RL. The performance evaluation was conducted using AD tasks and confirmed the ability of the developed algorithm to improve the learning efficiency compared to state-of-the-art human guidance-based RL. In addition, a human behavior modeling mechanism is proposed to reduce the burden on humans. Additionally, the authors of [66] investigate semi-autonomous UAVs to support emergency response scenarios. Humans and multiple UAVs must work together as a team to successfully execute a time-critical mission. Therefore, a meta-model is proposed to describe the interaction between humans and multiple UAVs. The meta-model provides a language to describe the roles of UAVs and humans and autonomous decisions. They have used two different scenarios involving multiple semi-autonomous UAVs in emergency operations to demonstrate the use of the meta-model by formally specifying human interactions and points of intervention.
HITL-DRL can faciliate UAV operation in complex environments and enable mission planning, further HITL-DRL address high-level decision-making problems in AD. The authors of [39] state that collision-free navigation for UAVs can facilitate data collection from IoT devices. However, traditional path planning methods are energy and computationally intensive, while DRL-based approaches tend to trap UAVs in complex environments. They propose a hybrid collision avoidance method for real-time UAV navigation in complex environments with unpredictable obstacles. They first develop a HITL-DRL training module for mapless obstacle avoidance and then set up a global planning module that generates some points as waypoint guidance. In addition, they have developed a novel algorithm for target updating that combines the HITL-DRL training module and the global planning module. In a similar vein, the authors of[46] develop an intelligent human–UAV interaction approach in real time based on ML using wearable gloves. The objective is to enable mission planning for controlling a UAV. ML-based methods process the signal from the gloves and classify appropriate commands to include in the human–UAV interaction process via the interface. The inclusion is according to the task scheduling algorithm to facilitate sequential and fast operation.
Table VII: Key HITL Techniques for Different Applications
@@
The combination of human expertise with RL algorithms ensures that AVs can handle complex and diverse environments effectively. However, AV bottleneck in handling large volumes of data, such as surveillance data and unlabeled data necessitates the use of AL to manage the situation.
III-C Key Findings and Insights
HITL-RL in AVs can be realized through reward shaping, human-initiated actions, and interactive learning to guide the AVs in critical situations. This can improve the efficiency of tasks and the learning ability and performance of the AVs. In addition, humans and AVs can interact and cooperate to solve tasks. In this regard, it should be noted that long-term monitoring and guidance can be physically and mentally demanding for human participants in the HITL-RL. It should be noted that current HITL-RL methods often require expert-level demonstrations to ensure data quality and optimize performance improvement. However, expert guidance is expensive for large-scale real-world applications. The main benefit of HITL-RL for AVs is to promote public trust and enhance safety.
Real-time human-UAV interactions using wearable gloves and ML based methods showcase innovative technology’s potential to enhance mission planning and control. Human and UAV interactions can be described using meta-models in emergency response scenarios to highlight the importance of formalizing human intervention points to ensure successful mission execution. HITL-RL frameworks are versatile and can be adapted to various applications, from AD to underwater exploration and emergency response. This versatility underlines the broad applicability and potential impact of HITL approaches in enhancing AVs across different domains. Incorporating mechanisms for human behavior modeling reduces the burden on human operators, making the system more efficient and user-friendly.
Developing a framework that learns reward functions from human driving data can effectively manage the learning process for AD, ensuring the RL agent optimizes complex goals. This approach includes trajectory sampling, offline preference learning, and RL integration to model human decision-making and obtain reward functions. Incorporating human prior experience into the reward function and safety constraints helps guide AVs in complex scenarios, such as mandatory lane changes, and limits unsafe exploration during training. Enhancing HITL-RL methods by having humans supervise, intervene, and demonstrate during the learning process not only improves the safety of AVs but also optimizes traffic flow efficiency in mixed-traffic environments. Creating a value-based HITL-DRL algorithm to address high-level decision-making problems, such as multi-objective lane-change decisions, improves training efficiency and adaptability. This approach values human policy over the undertrained DRL policy, allowing the RL agent to mimic human behaviors more efficiently. Enhancing RL performance by using human-guided learning objectives, prioritized human experience replay, and human intervention-based reward shaping improves the navigation capabilities of UGVs. Below, we discuss open research questions for HITL-RL.
III-C1 Designing AV Systems for Improved Decision-Making
A major open research question in the field of HITL-RL is how to design AV systems that improve decision-making over time through human interaction. To address this, we propose a human-centered design for DRL based on [68]. First, formulate a clear, testable hypothesis, such as ”DRL can optimize traffic signal timings to reduce city-wide congestion by 30%, ensuring it justifies human interaction and data collection. Second, align the DRL design with human needs for safety, convenience, and sustainability, engaging stakeholders like researchers, commuters, and policymakers to assess impacts and maintain open communication. Finally, set specific, measurable, attainable, relevant, and timely goals, such as achieving a 30% reduction in traffic congestion within three years, and define metrics for both optimization and human-centric success.
III-C2 Integration of Human Feedback in HITL-RL

### v2

@@ II-A1 Navigation and Data Collection
Curriculum learning can significantly improve collision avoidance in UAV swarms [57]by gradually increasing the complexity of the tasks that the UAVs have to solve during their training process. This approach ensures that the UAVs first master simpler tasks, such as avoiding static objects or other UAVs in a sparse environment, before moving on to more challenging scenarios, such as dynamic environments with multiple moving objects or denser UAV formations. A particular challenge in UAV swarms is collision avoidance, where the UAVs must avoid other UAVs as well as fixed and moving objects. Yan etal. [49] study the flocking and obstacle avoidance problem for multiple fixed-wing UAVs in cluttered environments. A novel curriculum-based multi-agent DRL is developed to learn the decentralized flocking and obstacle avoidance policy. In this context, the collision-avoiding flocking task is decomposed into multiple subtasks and the number of subtasks to be solved is gradually increased. In the developed approach, a novel multi-agent actor-critic DRL algorithm and two knowledge transfer mechanisms are developed for online learning and offline transfer, respectively. Similarly, Yan etal. [53] study collision-free flocking control problems for scalable and large-scale fixed-wing UAV swarms by integrating CL with multi-agent DRL. The problem considers both collision avoidance and the large population of UAVs in the swarm. The key idea is to divide the training process into several learning phases and gradually expand the scale of UAV swarms. In addition, Braathen etal. [58] investigate navigating UAVs in complex 3D environments and develop a solution using curriculum-based DRL. The human expert divides the training into ten different levels with an increasing range of possible numbers of obstacles in the environment. The leveraged curriculum exposes the UAV to simpler environments initially and progressively to more complex ones.
@@
Moreover, data collection [61] and UAV landing tasks require arranging the input. Seong etal. [59] address the collection of data from wireless sensor nodes to minimize the Age of Information during data collection, while considering the energetic sustainability of UAVs. They introduce Multi-Agent Reinforcement Learning (MARL) in combination with CL and an evolutionary strategy to improve the performance of traditional MARL setups as the number of agents increases. Their method promotes MARL scalability and integrates energy-efficient charging mechanisms that effectively improve system performance in large-scale deployments. Moreover, the collaborative autonomous landing of a quadrotor UAV on a moving Unmanned Ground Vehicle (UGV) is challenging due to the need for accurate real-time tracking of the UGV and the adjustment of the landing policy. Wang etal. of [60] propose the Landing Vision System to offer rapid localization and pose estimation of the UGV. Then, they design an Automatic CL approach in conjunction with the Twin Delayed Deterministic Policy Gradient (TD3) that breaks down UAV landing control into a series of simpler subtasks, and learn the landing tasks under different conditions of UGV motions and wind interference. Table III summarizes the above findings.
II-A2 Autonomous Driving
Self-driving tasks such as intersection management and motion planning benefit from CL to facilitate the learning tasks. Peng etal. [22] present a reward-driven automated curriculum proximal policy optimization (RD-ACPPO) approach for self-driving tasks at unsignalized intersections. Simply, an adaptive CL technique is presented to accommodate a progressively increasing difficulty of learning tasks in AD. Their approach allows for the reasonable allocation of the importance weights for different curricula during various RL training periods. The RL policy trained by the proposed framework attained the highest task success rate. In addition, Peiss etal. [62] develop a novel CL scheme for motion planning in AD that incorporates traffic rule and map information into a GNN-RL agent whose state is extended by map and traffic rule information. They decompose the complicated learning task by applying CL for GNN-RL agent. The curriculum’s sequence of sub-tasks gradually increases the level of difficulty of the motion planning problem for the agent. Each of their sub-tasks contains a set of rewards, including new rewards for temporal-logic-based traffic rules for speed, safety distance, and braking. Moreover, the authors of [50] investigate the challenge of learning AD behaviors at urban intersections with DRL. They combine CL with DRL to reduce the required training time and improve agent performance. They formed a set of tasks to consider and used an automated method to create the curriculum for the training process to avoid the burden of human handcrafting. Table IV summarizes the above findings.
@@ II-A2 Autonomous Driving
CL can facilitate learning and training in CARLA [65] driving simulator and target search for UAV swarms. Anzalone etal. [54] aim to learn an end-to-end urban driving policy for the CARLA driving simulator by combining the PPO algorithm with CL. In their approach, the RL phase is divided into several stages of increasing difficulty, so that the agent is guided to learn a better driving policy. The curriculum-based agent is evaluated on different metrics, cities, weather conditions, and traffic scenarios. Furthermore, Brunnbauer etal. [64] develop a multi-agent, scenario-based AD training and evaluation framework for the CARLA simulator. The results demonstrate the proposed approach is practical for multi-agent learning and can generate auto-curricula from scenario descriptions. In a similar vein, Xiao etal. [63] use UAV swarm for target search in disaster management scenarios. An adaptive embedded curriculum can adjust the difficulty of the task based on the success rate achieved in training and is used to enable UAVs to explore and navigate cluttered 3D environments with sparse rewards efficiently.
II-B Key Findings and Insights
@@
From the analysis in the previous subsection it became clear that CL, by systematically increasing the complexity of tasks, not only lays a solid foundation for the AVs learning process but also reduces training time and improves agent performance. This structured approach is crucial for preparing the vehicle to handle complex scenarios, where HITL-RL allows further fine-tuning the vehicle decision-making abilities.
In the sequel of our literature survey we highlight the following open issues:
II-B1 Merger with Transfer Learning
@@
II-B2 CL vs Hard Example Mining
When developing a curriculum for AVs, a common approach is to sort tasks from easy to hard, which is central to CL. However, this raises the question of whether this incremental approach is always the most effective strategy. An alternative approach, Hard Example Mining (HEM) [67], shifts the focus from a structured progression to selecting the most challenging examples for training. HEM operates under the assumption that harder examples, which are often identified by the model losses or gradient magnitude, provide more informative and valuable learning opportunities.
II-B3 Human Defined CL vs Human Data
@@
II-B4 Co-learning: a multi-agent approach to CL
@@ III Human-In-The-Loop Reinforcement Learning
HITL-RL significantly enhances the RL process by leveraging techniques like reward shaping, action injection, and interactive learning. These methods assist AVs in tasks such as lane-change, driving, obstacle avoidance, navigation, and overall performance enhancement, thereby improving safety and fostering public trust. The setup of HITL-RL for AVs is shown in Fig. 3. In this setup, human experts play a crucial role in monitoring the RL process. They can intervene by injecting actions when necessary and can enhance the system by providing supplementary rewards, known as human rewards. This integration ensures that the learning process remains safe, reliable, and aligned with human values and judgments.
III-A Literature review
III-A1 Reward Shaping
In reward shaping, the environmental reward is supplemented with human-defined rewards to guide learning. HITL can switch between multiple UAV reward functions in different situations and assist in constructing interactive reward functions. Humans can help to switch between different reward functions in different situations and replace the manual construction of the reward function.
Li etal. [41] develop a HITL-DRL framework for UAV navigation in large-scale 3D complex environments. This work allows the UAV to avoid obstacles automatically during flight. The role of HITL is to assist the UAV in obstacle avoidance by dynamically switching between multiple UAV reward functions in different situations. The combined approach significantly improves the success rate and navigation efficiency. In a similar vein, Karumanchi etal. [68] address the problem of autonomous UAV navigation and develop a practical solution based on Preference-based Reinforcement Learning (PbRL). In PbRL, the manual construction of the reward function is replaced with human feedback. A human is presented with a pair of trajectories followed by the RL agent and is asked to indicate a preference for one over the other. The PbRL algorithm then computes an optimal sequence of actions using only the set of preferences collected over different trajectories.
Table V: Key HITL-DRL Techniques with Reward Shaping
@@ III-A1 Reward Shaping
In the AD context, reward can be learned from human driving data and encoded with human prior experience. Zhuang etal. [69] address the intricate challenges in designing appropriate reward functions to manage optimizing complex goals in RL. They develop a framework that learns the reward function from human driving data to help manage the learning process. Trajectory sample, offline preference learning, and RL are the comprising components, where feasible trajectories are generated by sampling end targets from a reachable state space, then a transformer network that is used to model the human driving decision-making process, and obtaining a reward function, is trained by a new offline preference learning where generated feasible trajectories are compared with human driving trajectories. Finally, the final driving decision network is obtained by incorporating the derived reward function in the RL framework. In another related effort, Huang etal. [70] aim to facilitate AVs mandatory lane-change scenarios in complex environments. In this context, they develop a HITL-RL method based on Twin Delayed Deep Deterministic policy gradient (TD3) for changing lane decisions. They encode human prior experience into reward function and safety constraints in an offline manner. Also, human experts guide the agents online to limit unsafe exploration during the training process and provide a demonstration in complex scenarios. Table V summarizes the above findings.
III-A2 Action Injection
@@
Introducing human guidance into the training loop of DRL algorithms enable AV to better handle different situations. Further, human intervention can make the UAV swarm safe. Wu etal. [42] state that the limited intelligence and capabilities of DRL-based strategies in AV applications prevent them from handling various situations and replacing human drivers. Introducing human guidance into the training loop of DRL algorithms can improve the situation. In the proposed HITL-DRL, human guidance is used in real-time to improve the agent’s learning ability and performance. Specifically, the human participant observes the agent’s real-time training process and assists when necessary. The provided action of the human guidance is prioritized as the output action of the agent to interact with the environment and replaces the original action of the DRL policy. The actor and critic are updated according to the human guidance and experience of the DRL agent. The actor-network can learn from both human guidance through imitation learning and the experience of the interactions through RL.
Table VI: Key HITL-DRL Techniques with Action Injection
@@
The critic network can evaluate the values of the agent’s actions as well as the actions of human guidance. Similarly, Li etal. [71] develop a HITL-DRL framework for UAV swarm navigation in a large-scale 3D complex urban environment. This framework enables human intervention in the flight of the UAV swarm when necessary to correct dangerous actions. Additionally, the Actor-Critic structure of DDPG has been modified to identify and learn from positive human experiences. Human feedback can enhance UAV obstacle avoidance performance by having the human acting as a mentor to the AI agent of AD. Lee etal. [72] investigate autonomous navigation in challenging environments with the capability to avoid obstacles. The UAV navigates toward its target using the DDPG model, which can handle unexpected situations. HITL feedback is integrated with the DDPG model to enhance obstacle avoidance performance. This feedback is crucial when the pre-trained DDPG model encounters environments it is not fully adapted to, and gives the system the flexibility to handle unforeseen scenarios effectively and make real-time adjustments. For the HCI feedback, the first-person shooter concept is employed, where feedback is provided using keyboard directional keys. Through this HCI-based HITL feedback, real-time control of the UAV flight movement is achieved.
Huang etal. [73] aim to improve the safety of AD technology, ensuring traffic flow efficiency in mixed-traffic environments. They develop an enhanced HITL-RL for driving policy learning, where the human serves as a mentor to the AI agent, supervising, intervening, and demonstrating in its learning process, and integrated with DRL to guarantee the safety of the AVs, but also optimize the traffic flow efficiency. Table VI summarizes the above findings.
III-A3 Interactive Learning
@@
The agent and human trainer can interact and cooperate to solve tasks. For example, human can supplement environmental reward and guide learning objectives to improve learning efficiency and RL performance. Zhang etal. [43] aim to explore the oceans using autonomous underwater vehicles (AUV). RL can improve its autonomy but still has practical problems due to sparse rewards and low learning efficiency. They propose a deep interactive RL method for AUV path tracking by combining the advantages of DRL and interactive RL. Since the human trainer cannot give human rewards to the AUV in the operational condition and the AUV needs to adapt to a changing environment when it is in the ocean, a DRL method that simultaneously learns from human rewards and environmental rewards is proposed. Wu etal. [74] improve RL performance in simulation and practice for the navigation task of UGVs using HITL-RL. The framework improves human interventions with a series of mechanisms to provide demonstrations and improve RL’s capabilities. The mechanisms comprise human-guided learning objectives, prioritized human experience replay, and human intervention-based reward shaping.
@@
Human behavior modeling mechanisms can alleviate the workload on humans, further using meta-model provides a language to describe the roles of UAVs and humans and autonomous decisions. Wu etal. [45] aim to improve the learning performance of RL by incorporating human guidance. In this regard, they develop a human guidance-based RL framework along with a novel mechanism for leveraging the experience of human guidance. A TD3-based algorithm was developed to improve algorithmic capabilities in the context of HITL-RL. The performance evaluation was conducted using AD tasks and confirmed the ability of the developed algorithm to improve the learning efficiency compared to state-of-the-art human guidance-based RL. In addition, a human behavior modeling mechanism is proposed to reduce the burden on humans. Additionally, Agrawal etal. [75] investigate semi-autonomous UAVs to support emergency response scenarios. Humans and multiple UAVs must work together as a team to successfully execute a time-critical mission. Therefore, a meta-model is proposed to describe the interaction between humans and multiple UAVs. The meta-model provides a language to describe the roles of UAVs and humans and autonomous decisions. They have used two different scenarios involving multiple semi-autonomous UAVs in emergency operations to demonstrate the use of the meta-model by formally specifying human interactions and points of intervention.
Table VII: Key HITL Techniques with Interactive Learning
@@
HITL-DRL can faciliate UAV operation in complex environments and enable mission planning, further HITL-DRL address high-level decision-making problems in AD. Zhang etal. [44] state that collision-free navigation for UAVs can facilitate data collection from IoT devices. However, traditional path planning methods are energy and computationally intensive, while DRL-based approaches tend to trap UAVs in complex environments. They propose a hybrid collision avoidance method for real-time UAV navigation in complex environments with unpredictable obstacles. They first develop a HITL-DRL training module for mapless obstacle avoidance and then set up a global planning module that generates some points as waypoint guidance. In addition, they have developed a novel algorithm for target updating that combines the HITL-DRL training module and the global planning module. In a similar vein, Muezzino ¨ glu etal. [51] develop an intelligent human–UAV interaction approach in real time based on ML using wearable gloves. The objective is to enable mission planning for controlling a UAV. ML-based methods process the signal from the gloves and classify appropriate commands to include in the human–UAV interaction process via the interface. The inclusion is according to the task scheduling algorithm to facilitate sequential and fast operation.
@@
III-B Key Findings and Insights
From the discussion in the previous subsection we could see that the combination of human expertise with RL algorithms enables AVs to handle complex and diverse environments effectively. The potential AV bottleneck in handling large volumes of data, such as surveillance data and unlabeled data, can be addressed using AL to manage the situation. Below, we discuss open research questions concerning HITL-RL:
III-B1 Designing AV Systems for Improved Decision-Making
A major open research question in the field of HITL-RL is how to design AV systems that improve decision-making over time through human interaction. To address this, a possible approach is to follow a human-centered design for DRL [77]. First, a clear and testable hypothesis is formulated, e.g., DRL can optimize traffic signal timings to reduce city-wide congestion by 30%, ensuring it justifies human interaction and data collection. Second, the DRL design is aligned with human needs for safety, convenience, and sustainability, engaging stakeholders like researchers, commuters, and policymakers to assess impacts and maintain open communication. Finally, specific measurable, attainable, relevant, and timely goals are set, such as achieving a 30% reduction in traffic congestion within three years, and metrics are defined for both optimization and human-centric success.
III-B2 Integration of Human Feedback in HITL-RL